OpenStudy (anonymous):

convert to decimal notation...8.92x10^7=89,000,000?

OpenStudy (lilg132):

8.92 x10^1 = 89.2 x10^2 = 892 x10^3 = 8,920 x10^4 = 89,200 x10^5 = 892,000 x10^6 = 8,920,000 x10^7 = 89,200,000 dont forget to click good answer :)

OpenStudy (anonymous):

ugh i forgot the 2

OpenStudy (anonymous):

wait its 8.92 so wouldnt it be 892,000,000

OpenStudy (lilg132):

no lol how did you get that?

OpenStudy (lilg132):

remember 8.92 x10^1 is 89.2

OpenStudy (lilg132):

so just move it 6 decimal places to the left after you have that

OpenStudy (lilg132):

sorry move it 6 places to the left

OpenStudy (lilg132):

which will give you 89200000

OpenStudy (lilg132):

89,200,000

OpenStudy (lilg132):

8.92 turning into 892,000,000 would mean 8.92 x10^8

OpenStudy (anonymous):

its to the 7th power

OpenStudy (lilg132):

so it will be 89,200,000

OpenStudy (lilg132):

if you have 8.92 and it says what is the number if it is x10^7 that will mean you move all the numbers to the left 7 times.

OpenStudy (lilg132):

is that easier to understnad?